OpenStudy (rational):

[World cup fun problem!] \(\large 7\) countries are participating in a tournament. In the first round, each country plays a match against every other country and the lowest scoring \(3\) teams will be eliminated. Points are awarded as below : ``` win : 2 draw : 1 loss : 0 ``` Find 1) Minimum points required by a team that guarantee surviving the first round. 2) All the possible permutations of points possible after first round.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|c|c|c|c|c|} \hline \text{teams} & \text{ A} & \text{B}& \text{ C} & \text{D}& \text{E} & \text{F}& \text{G} &\text{total}\\ \hline \text{A} & -& 0& 1 & 1& 1 & 1& 1 & 5\\ \hline \text{B} & 2& -&0 & 0& 1 & 1& 1& 5 \\ \hline \text{C} & 1& 2&- & 1& 0 & 0& 0&4 \\ \hline \text{D} & 1& 2&1 & -& 1 & 1& 1&7 \\ \hline \text{E} & 1& 1&2 & 1& -& 1& 1& 7\\ \hline \text{F} & 1& 1&2 & 1& 1&-& 1& 7\\ \hline \text{G} & 1& 1&2 & 1& 1&1& - &7\\ \hline \end{array} \end{align}}\)

OpenStudy (mathmath333):

from this table teams \(\large A,B\quad \text{and}\quad C\) are eliminated and the minimum point seems \(\large 7\)

OpenStudy (rational):

your table actually shows that \(6\) points will guarantee survival because the 3 eliminated teams have points less than \(6\)

OpenStudy (rational):

look at the last column for points http://gyazo.com/f9a654a36f3c59f2ed76858e55ebcb1e If your table is correct, it seems india and south africa have cleared the first round ?

OpenStudy (rational):

the question then is : can a team get eliminated even after scoring 6 points ?

OpenStudy (mathmath333):

yes it is \(6\) but i just selected an arbitrary case, where the 4 teams had \(7\) points

OpenStudy (rational):

then why india and south africa are not celebrating yet lol

OpenStudy (mathmath333):

i honestly don't watch cricket , and they have already celebrated i think, they are in last six , so way ahead to go, do or die every match

OpenStudy (rational):

Okay but it seems india and south africa have already scored the minimum required points(6). So they will survive even if they lose all remaining matches

OpenStudy (mathmath333):

yes they will survive guaranteed.

OpenStudy (rational):

(again thats according to your table) I haven't seen any news channels mentioning about these countries clearing the first round yet..

OpenStudy (mathmath333):

but the run rate matters in case they tie in top six , so every run\win counts

OpenStudy (mathmath333):

i have read this in newspaper i think

OpenStudy (rational):

is there a chance for more than four teams to score >= 6 points ?

OpenStudy (mathmath333):

in world cup it is not single round robin in first phase i think , but in your question it is right

OpenStudy (mathmath333):

the maximum matches a team can play in first phase is \(4\) , i think

OpenStudy (rational):

The tournament will feature 14 teams, the same number as the 2011 World Cup. The format is the same as the 2011 edition: 14 teams will take part in the initial stages, divided into two groups of seven; the seven teams play each other once before the top four teams from each group qualify for the quarter-finals.

OpenStudy (rational):

The format ensures that each team gets to play a `minimum of six matches` even if they exit in the group stage.

OpenStudy (mathmath333):

oh i see , it is single round robin

OpenStudy (rational):

yes

OpenStudy (rational):

so it all boils down to proving below : There is no chance for more than four teams to score \(\ge 6\) points

OpenStudy (rational):

that will prove 6 is the minimum required points for guaranteeing survival

OpenStudy (mathmath333):

yep six!

OpenStudy (mathmath333):

chess tournament has also the similar scoring system

OpenStudy (mathmath333):

but in case of tie , the two players play a duel match

OpenStudy (mathmath333):

*tie for the first place

OpenStudy (rational):

yeah rapid round

OpenStudy (rational):

so how do we prove 6 is the minimum ?

OpenStudy (mathmath333):

draw a table where 4 teams have 6 points

OpenStudy (rational):

is this possible ? \[\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|c|c|c|c|c|} \hline \text{teams} & \text{ A} & \text{B}& \text{ C} & \text{D}& \text{E} & \text{F}& \text{G} &\text{total}\\ \hline \text{A} & -& 0& 1 & 1& 1 & 1& 1 & 5\\ \hline \text{B} & 2& -&2 & 0& 1 & 1& 1& 7 \\ \hline \text{C} & 1& 0&- & 1& 0 & 0& 0&2 \\ \hline \text{D} & 1& 2&1 & -& 1 & 1& 1&7 \\ \hline \text{E} & 1& 1&2 & 1& -& 1& 1& 7\\ \hline \text{F} & 1& 1&2 & 1& 1&-& 1& 7\\ \hline \text{G} & 1& 1&2 & 1& 1&1& - &7\\ \hline \end{array} \end{align}}\]

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \begin{array}{|c|c|c|c|c|c|c|c|c|} \hline \text{teams} & \text{ A} & \text{B}& \text{ C} & \text{D}& \text{E} & \text{F}& \text{G} &\text{total}\\ \hline \text{A} & -& 0& 1 & 1& 1 & 1& 1 &\color{red}{ 5}\\ \hline \text{B} & 2& -&0 & 0& 1 & 1& 1& \color{red}{ 5} \\ \hline \text{C} & 1& 2&- & 2& 0 & 0& 0&\color{red}{ 5} \\ \hline \text{D} & 1& 2&0 & -& 1 & 1& 1&\bf\color{black}{ 6} \\ \hline \text{E} & 1& 1&2 & 1& -& 1& 1& \color{green}{ 7}\\ \hline \text{F} & 1& 1&2 & 1& 1&-& 1& \color{green}{ 7}\\ \hline \text{G} & 1& 1&2 & 1& 1&1& - &\color{green }{ 7}\\ \hline \end{array} \end{align}}\) this case shows the minimum is \(\bf \Large 6\)

OpenStudy (ikram002p):

But why u didn't consider each game in first round as only lose or win game ? And how are the countrties gonna play like how many games in.first round allowed ?

OpenStudy (ikram002p):

Oh wait my b.ad huh got it now

OpenStudy (mathmath333):

it can be possible to consider ,where a country either losses or win , but the their are many combinations but i chose maximum draws to make it simpler

OpenStudy (mathmath333):

there are total of seven countries and each country will play total of 6 games